# Accreditation Against Limited Adversarial Noise

### v2

@@ II.2 Problem Setting: Preliminaries and Definitions
For any given redaction class and any given PSCLβ either:
$\mathbb{1.}$ Every CPTP list in the PSCLβ fits every circuit in the redaction class
@@
$\mathbb{2.}$ No CPTP list in the PSCLβ fits any circuit in the redaction class.
I refer to the first of these possibilities as the redaction class and the PSCLβ matching, but also allow the term to apply if the redaction class is replaced by any subset of itself.
@@
As mentioned in Def. 7 if one CPTP list in a given PSCLβ fits a specific circuit then the entire PSCLβ fits that circuit.
The lemma follows from combining these two facts. ∎
@@ II.3 Problem Setting Presentation
I can now present the protocol by which sets of circuits, in my problem setting, are executed, in Protocol 1.
@@
    Alice provides a set of circuits to execute, $\mathcal{S}$, all within the same redaction class.
@@
    Bob receives the full details of $\mathcal{S}$, with all single-qubit gates in every element of $\mathcal{S}$ redacted.
@@
    Bob chooses a SPSCLβ that matches $\mathcal{S}$, denoted $\Xi$, where $\beta\in\mathbb{R}^{+}$ is known to both Alice and Bob.
@@
      Bob chooses a CPTP list, $\xi$, from $\Xi$.
@@
      $\mathcal{C}$ is executed with $\xi$ determining the error occurring during the execution.
@@
      Alice and Bob are both given the measurement outcomes of the execution of $\mathcal{C}$.
@@
Alice is free to initiate Protocol 1 whenever and as often as she likes; Bob is free to act how he likes to achieve his stated malicious aims but must execute Protocol 1 as prescribed - but is free to act nefariously within the permitted bounds - when it is invoked. While the bulk of this paper focuses on constructing the accreditation protocol protocol, without much regard for Alice, it culminates in Theorem 2 showing that Alice’s problem is solved by the protocol presented herein.
II.4 Physical Origins of the Error Model / Problem Setting
The two main limitations on the adversarial error / Bob requiring justification are:
@@ II.4 Physical Origins of the Error Model / Problem Setting
    The single-qubit gates are redacted when the circuits are shown to Bob (in step 2 of Protocol 1).
@@
    Within a single use of Protocol 1, Bob is restricted to only causing error describable by CPTP lists with a single SPSCLβ, for a known $\beta\in\mathbb{R}^{+}$.
The first limitation corresponds to single-qubit gates experiencing gate-independent error. This is a standard assumption the the pre-existing accreditation protocols (in Refs. [22, 23, 24]) and follows from single-qubit gates typically being the least error-prone components of a quantum computer [28, 29] (which has held true over time [30, 31], and can be most clearly seen in Ref [32, Fig.5]): the error is so small that the error in different single-qubit gates does not differ much. This is a “standard [assumption] in the literature on noise characterisation and mitigation” [33] and has seen extensive use in theoretical work [3, 34, 35, 36, 37, 38, 39, 40, 41, 42, 33].
The second limitation is a weakening of the IID assumption in previous accreditation protocols. The intuition behind it is that the same hardware executing very similar circuits in quick succession - as happens in Protocol 1 will experience similar error in each circuit as:
@@
    The external factors affecting error have to be fluctuating over a very specific time-span - slow enough to not fluctuate much within the execution of circuits but fast enough to change between executions of a circuit - to cause significant differences in the error in different executions executed in quick succession.
@@
Although no paper has, to my knowledge, sought to directly validate this limitation, it can also be justified experimentally:
@@
    Ref. [23, Figure. 5] ran the accreditation protocol - developed therein - many times, producing a probability of error in the traps for each use of the protocol. The set of error probabilities generated by this can be seen to vary little across many uses of the protocol.
@@
    Ref. [43, Fig. 6, Fig. 7, and Fig. 8] investigated the error rates in single-qubit and two-qubit gates. It found these error rates are very rarely far from their average. Note that the variation shown in this paper is over a much longer time-span (days) than this limitation requires (seconds).
@@
    Ref. [32, Fig. 9] examined NISQ devices and - across many devices - plotted the error rates of different qubits in each device. Fig. 9 shows the error bars on the error rate to be - with some notable exceptions - small fractions of the error rate and varying much more across qubits than for a fixed qubit over time.
  4. 4.
    Ref. [44, Fig. 1(b)] looked at the error rates of two-qubit gates on differing pairs of qubits, it shows that - again, over a time-span of days - for a specific pair of qubits, the days when the error rate deviates far from its average are rare, with more extreme deviations being rarer.
Finally, this limitation can be seen as an aim of quantum computer hardware engineering: as quantum computers improve and their actual outputs approach the errorless outputs, with decreasing variance, the typical difference - by any measure - between the error contained in two executions of similar circuits will tend to zero.
To my knowledge, I am the first to explicitly state and use this second limitation. However, I note that it is already implicitly accepted in the community by the acceptance of randomized benchmarking [3, 4, 5, 6] as a meaningful measure: if even the same circuit repeated multiple times, in very quick succession, produces wildly varying probabilities of error then the variation of error probabilities implies that previous measures mean almost nothing for future computations or the quality of a quantum device over a meaningful time-span. This is not the case and Ref. [44, Fig. 6] shows that, over a span of days, the results of randomized benchmarking do not vary over a very large range.
There is, however, another, subtle, restriction on Bob: he controls the CPTP maps acting on all executions, but does not control any stochastic elements of the implementation of the CPTP map. I.e. he may specify that a specific CPTP map in the execution applies stochastic Pauli error, such as a dephasing channel, but he does not decide if this results in a Pauli $\hat{Z}$ gate being applied. Furthermore, Bob is not informed of the outcomes of these stochastic processes - in spite of choosing which stochastic processes occur. This represents that the environment - where any memory of past error that the noise represented by Bob has must be - is not, consistently or over long time-spans, protected from noise acting on it and destroying any information/memory it contains, so any hostile intelligence that lurks in the noise must have a comparatively short and small memory  [^8] Quantum hardware engineers have expended much effort to keep the the information inside quantum computers intact, so there is no reason to assume the environment can - with non-negligible probability - store significant amounts of information, for relevant timespans, without any engineering effort to enable it to do so..
III Accreditation Protocol
Sec. III is dedicated to resolving the problem Alice faces in the problem setting established in Sec. II.3. This is equivalent to upgrading an accreditation protocol that assumed error is CPTP and IID (and that single-qubit gates suffer only gate-independent error) to one that works with the error assumptions in Sec. II.3.
III.1 Trap and Target Circuits
In this paper, I do not propose to develop new trap circuits or target circuits. In fact, I would prefer to make minimal changes to the trap and target circuits in Ref. [23]. I will also not regurgitate the exact designs of the trap and target circuits in Ref. [23] and instead - throughout this paper - will assume that I have two efficient classical algorithms, $P_{\textit{targ}}$ and $P_{\textit{trap}}$, that, if given any quantum circuit as input, return the corresponding target and a random [^9] I.e. $P_{\textit{trap}}$ will choose the random gates for the Pauli twirl and error detection and so will return a slightly different circuit each time. matching trap, respectively, of the protocol in Ref. [23].
I briefly note the important features of these trap and target circuits:
  1. 1.
    In target circuits and trap circuits all error occurring is twirled (via Pauli twirls or some other twirl) to stochastic Pauli error.
  2. 2.
    Each execution of a trap acts as a Bernoulli trial, with the success probability - for that trial - being the probability of error occurring in that trap execution.
  3. 3.
  4. 4.
    Target circuits and trap circuits differ only in their single-qubit gates.
The above assumptions have mentioned twirling CPTP error to stochastic Pauli error, which I define formally in Def. 8. This is an important step in all accreditation protocols, as it reduces the error to a known, more easily quantified form.
@@ III.1 Trap and Target Circuits
CPTP error within a quantum circuit is said to be twirled to stochastic Pauli error [45, 46, 47] if, via the addition of only single-qubit Pauli gates to the circuit, the error is effectively transformed to stochastic Pauli error, without otherwise affecting the outputs of the circuit. Likewise, gates are said to be twirled if any error occurring in them is twirled.
As herein noise is now considered to be adversarial, the accreditation protocol of this paper does require a slight modification of the trap circuits and target circuits: Ref. [23] did not consider “hiding” the measurement outcomes of a circuit as there was no adversary to “hide” them from (according to its error model). With adversarial noise (i.e. the problem setting of Sec. II.3), this becomes necessary as otherwise Bob may be able to identify traps and targets based on these outcomes, or base future error on the measurement outcomes of previous circuits. The required hiding / encrypting of measurement outcomes is achieved via Algorithm 2.
@@
$\bullet$ A circuit, $\mathcal{C}$, to generate traps or a target for.
$\bullet$ Two algorithms, $P_{\textit{trap}}$ and $P_{\textit{targ}}$, as detailed above.
$\bullet$ A Boolean, labelled isTarget, denoting if a trap or target is to be generated.
@@
Algorithm 2 Generating Traps and Targets with Hidden Outputs
I note that the required (by Algorithm 2) single-qubit unitaries, $\mathcal{U}_{M}$, will exist for any single-qubit measurement and that the outcome of the circuit - if the single-qubit gates of the circuit are not known, as is the case for Bob as they are redacted for him - is irretrievable from the measurement outcomes without the key. This hiding of the circuit outputs is information-theoretically secure (i.e. has perfect security) [48], assuming the single-qubit gates are redacted, and comparable to the one-time pads in universal blind quantum computing [49]. Henceforth, I will denote $P_{\textit{targ}}$ and $P_{\textit{targ}}$, with the changes presented above to make their outputs hidden, by $P_{\textit{targ}}^{\prime}$ and $P_{\textit{targ}}^{\prime}$ respectively.
@@ III.2 Presentation of the Upgraded Accreditation Protocol
With trap and target circuits established, I can consider the full accreditation protocol to be developed. In line with Sec. III.1, for Sec. III.2, traps and targets will be treated as blackboxes and I will only refer to their construction as being performed by $P_{\textit{targ}}^{\prime}$ and $P_{\textit{targ}}^{\prime}$. Their only relevant properties will be that traps detect any error with probability $k\in[0,1]$, all error in a trap or target simulation is reduced to effective Pauli error, and Bob cannot tell the difference between trap and target simulations due to them only differing in their single-qubit gates.
However, before the full protocol can be presented, in Protocol 3, I must establish the statistical foundations of the new protocol - in Sec. III.2.1 - and the core mechanics of the protocol - in Sec. III.2.2. These statistical methods will be used to evaulate the $P_{0}$ that defines Bob’s particular choice of a SPSCLβ using multiple trap executions (still all within a single use of Protocol 1).

### v3

@@ II.2.3 Relating CPTP Lists and Redaction Classes
For any given redaction class and any given SPSCLβ $\mathbb{either}$:
$\mathbb{1.}$ Every CPTP list in the SPSCLβ fits every circuit in the redaction class
@@
$\mathbb{2.}$ No CPTP list in the SPSCLβ fits any circuit in the redaction class.
I refer to the first of the above options as the redaction class and the SPSCLβ matching (but also allow the term to apply if the redaction class is replaced by any subset of itself).
@@
As in Def. 12 if one CPTP list in a given SPSCLβ fits a specific circuit then every CPTP list in that SPSCLβ fits that circuit.
Lemma 2 follows from combining these two facts. ∎
@@ II.3 Problem Setting Presentation
I can now present the protocol by which sets of circuits, in my problem setting, are executed, in Protocol 1. Both Alice and Bob are required to participate in Protocol 1 and both must abide by all requirements of it. These requirements are enforced by another character, Robert, who actually implements the quantum computations – in Protocol 1 – according to Alice and Bob’s combined instructions. However, Protocol 1 still leaves room for Bob to attempt to trick Alice.
@@
    Alice provides a set of circuits to execute, $\mathcal{S}$, all within the same redaction class, to Robert.
@@
    Bob receives the full details of $\mathcal{S}$, from Robert, but with every circuit in $\mathcal{S}$ redacted.
@@
    Bob chooses an SPSCLβ, denoted $\Xi$, that matches $\mathcal{S}$ and tells Robert it, where $\beta\in\mathbb{R}^{+}$ is known to both Alice and Bob in advance.
@@
      Bob chooses a CPTP list, $\xi$, from $\Xi$ and gives it to Robert.
@@
      $\mathcal{C}$ is executed, by Robert, with $\xi$ determining the error occurring during the execution.
@@
      Robert gives both Alice and Bob the measurement outcomes of the execution of $\mathcal{C}$.
@@
Alice is free to initiate Protocol 1 whenever and as often as she likes; Bob is free to act how he likes to achieve his stated malicious aims but must execute Protocol 1 as prescribed – but is free to act nefariously within the permitted bounds – when it is invoked. Robert checks Bob’s nefarious choices are within the rules of Protocol 1 and executes circuits.
The aims of Alice and Bob remain the same as in the standard cryptographic setting and as discussed in Sec. II.1: Alice is attempting to get Robert (via Protocol 1) to perform a quantum computation and provide her with the true results, while Bob is trying to trick Alice by adding error to Alice’s requested circuit (when Robert executes it in Protocol 1) so that Robert provides Alice with the results of an incorrect computation and Alice believes that the computation was performed without error. I.e. that the results Alice receives are the outputs of the computation she wanted to be performed when they are not (due to the error Bob adds). However, in trying to achieve these aims, Alice and Bob can only communicate via Protocol 1 with Robert. Robert’s aims are entirely neutral, he has no preference on if Alice or Bob successfully achieves their aim. He just wants to do his duty and perform Protocol 1 correctly.
The problem setting of this paper is summarized in Table 2.
Alice Bob Robert
Aim Obtain the results of a specific computation Alice to accept incorrect results Implement Protocol 1 as requested
Computational Capabilities Polynomial time classical computation Unbounded sampBQP computations
Additional Capabilities May initiate Protocol 1 at any time None None
Allowed Communication None, aside from via Protocol 1 None, aside from via Protocol 1 None, aside from as in Protocol 1
Table 2: Table summarizing the problem setting of this paper (i.e. the adapted cryptographic setting) in terms of the aims, computational abilities, additional capabilities, and allowed communication of its only characters: Alice, Bob, and Robert.
While the bulk of the rest of this paper focuses on constructing the accreditation protocol, without much regard for Alice, it culminates in Theorem 1 showing that Alice’s problem is solved by the protocol presented herein. I.e. in the problem setting from this section, the protocol presented in this paper allows Alice to get Bob to perform a quantum computation for her and have confidence in the results she receives.
II.4 Physical Justification of the Problem Setting
The two main limitations on the adversarial error / Bob, both following from Protocol 1, requiring justification are:
@@ II.4 Physical Justification of the Problem Setting
    The single-qubit gates are redacted when the circuits in $\mathcal{S}$ are shown to Bob (in step 2 of Protocol 1).
@@
    For all circuit executions within a single use of Protocol 1, the different CPTP lists Bob uses to determine the error in each execution are all within a single SPSCLβ, for a known $\beta\in\mathbb{R}^{+}$.
Both of these limitations are applied (in the case of the first) or enforced (in the case of the second) by Robert.
The first limitation corresponds to single-qubit gates experiencing gate-independent error. This is a standard assumption the the pre-existing accreditation protocols (in Refs. [22, 23, 24]) and follows from single-qubit gates typically being the least error-prone components of a quantum computer [30, 31] (which has held true over time [32, 33], and can be most clearly seen in Ref [34, Fig. 5]): the error is so small that the error in different single-qubit gates does not differ much. This is a “standard [assumption] in the literature on noise characterisation and mitigation” [35] and has seen extensive use in theoretical work [3, 36, 37, 38, 39, 40, 41, 42, 43, 44, 35].
The second limitation is a weakening of the IID assumption in previous accreditation protocols. The intuition behind it is that the same hardware executing very similar circuits in quick succession – as happens in Protocol 1 – will experience similar error in each circuit execution as:
@@
Although no paper has, to my knowledge, sought to directly validate this limitation, it can be justified experimentally:
  1. 1.
    Ref. [23, Figure. 5] ran the accreditation protocol developed therein many times, producing a probability of error in the trap circuit executions for each use of the protocol. The set of error probabilities generated by this can be seen to vary little across many uses of the protocol.
  2. 2.
    Ref. [45, Fig. 6, Fig. 7, and Fig. 8] investigated the error rates in single-qubit and two-qubit gates. It found these error rates are very rarely far from their average. Note that the variation shown in this paper is over a much longer time-span (days) than this limitation requires (seconds).
@@
    Ref. [34, Fig. 9] examined NISQ computers and plotted the error rates of different qubits in each device. Fig. 9 shows the error bars on the error rate to be – with some notable exceptions – small fractions of the error rate and varying much more across qubits than for a fixed qubit over time.
  4. 4.
    Ref. [46, Fig. 1(b)] looked at the error rates of two-qubit gates on differing pairs of qubits, it shows that (again, over a time-span of days) for a specific pair of qubits, the days when the error rate deviates far from its average are rare, with more extreme deviations being rarer.
Finally, this limitation can be seen as an aim of quantum computer hardware engineering: as quantum computers improve and their actual outputs approach the errorless outputs, with decreasing variance, the typical difference – by any measure – between the error contained in two executions of similar circuits will tend to zero.
To my knowledge, I am the first to explicitly state this second limitation. However, I note that it is already implicitly accepted in the community by the acceptance of randomized benchmarking [3, 4, 5, 6] as a meaningful measure: if even the same circuit repeated multiple times, in very quick succession, produces wildly varying probabilities of error then the variation of error probabilities implies that previous measures mean almost nothing for future computations or the quality of a quantum device over a meaningful time-span. This is not the case and Ref. [46, Fig. 6] shows that, over a span of days, the results of randomized benchmarking do not vary over a very large range.
III Adversarial Accreditation Protocol
Sec. III is dedicated to resolving the problem Alice faces in the problem setting established in Sec. II.3. This is equivalent to upgrading the accreditation protocol in Ref. [23] that assumed error is CPTP and IID (and that single-qubit gates suffer only gate-independent error) to one that works in the problem setting described in Sec. II.3.
Sec. III begins – in Sec. III.1 – with a presentation of the trap and target circuits I intend to use in the new accreditation protocol, which are very similar to those in Ref. [23]. As a trap-based verification protocol, the accreditation protocol presented in Sec. III needs these trap circuit executions to be executed alongside the target circuit (in the same single use of Protocol 1 and hence experiencing comparable error by the assumptions implicit in the problem setting defined in Sec. II) and give a measure of the quality of the execution of the target circuit. The usage of these trap and target circuits to produce an accreditation protocol is them detailed in Sec. III.2 (and more formally presented in Algorithm 3).
III.1 Trap and Target Circuits
In this paper, I do not propose to develop new trap circuits or target circuits. In fact, I would prefer to make minimal changes to the trap and target circuits in Ref. [23]. I will also not regurgitate the exact designs of the trap and target circuits in Ref. [23] and instead throughout this paper will assume that I have two efficient classical algorithms, $P_{\textit{targ}}$ and $P_{\textit{trap}}$, that, if given any quantum circuit as input, return a random [^10] I.e. $P_{\textit{trap}}$ and $P_{\textit{targ}}$ will choose random gates used to apply the Pauli twirls and probabilistic error detection (via Hadamard gates), so will return a slightly different circuit each time. corresponding target and matching trap, respectively, of the protocol in Ref. [23]. I briefly note the important features of these trap and target circuits (that will be inherited wherever I use trap or target circuits herein):
@@ III.1 Trap and Target Circuits
    In target circuits and trap circuits all error occurring is twirled (via Pauli twirls) to stochastic Pauli error and is thereafter considered as such.
@@
    Target circuits and trap circuits differ only in their single-qubit gates so they are all in the same redaction class.
The above assumptions have mentioned twirling CPTP error to stochastic Pauli error, which I define formally in Def. 13. This is an important step in all accreditation protocols, as it reduces the error to a known, more easily quantified form.
@@
CPTP error within a quantum circuit is said to be twirled to stochastic Pauli error [47, 48, 49] if, via the addition of only single-qubit Pauli gates to the circuit, the error is effectively transformed to stochastic Pauli error, without otherwise affecting the outputs of the circuit (e.g. it does not affect the outputs of the errorless case). Likewise, gates are said to be twirled if any error occurring in them is twirled to stochastic Pauli error.
As herein noise is considered to be adversarial, the accreditation protocol of this paper does require a single slight modification of the trap circuits and target circuits: Ref. [23] did not consider “hiding” the measurement outcomes of a circuit as there was no adversary to “hide” them from (according to its error model). With adversarial noise (i.e. the problem setting of Sec. II.3), this becomes necessary as otherwise Bob may be able to identify which circuits are trap and target circuits, respectively, based on these outcomes, or base future error on the measurement outcomes of previous circuits (removing the independence of the error in different circuits). The required hiding / encrypting of measurement outcomes is achieved via Algorithm 2 which; using the classical algorithms for generating the trap and target circuit of Ref. [23], $P_{\textit{trap}}$ and $P_{\textit{targ}}$ respectively; acts as a classical algorithm to generate trap and target circuits similar to those of Ref. [23] but with the outputs quantum-securely encrypted and completely unrecoverable without the key.
@@
$\bullet$ A circuit, $\mathcal{C}$, to generate trap circuits or a target circuit for.
$\bullet$ Two algorithms, $P_{\textit{trap}}$ and $P_{\textit{targ}}$, that generate the required trap and target circuits, respectively.
$\bullet$ A Boolean, labeled isTarget, denoting if a trap or target is to be generated.
@@
Algorithm 2 Generating Trap and Target Circuits with Hidden Outputs
I note that the required (by Algorithm 2) single-qubit unitaries, $\mathcal{U}_{M}$, will exist for any single-qubit measurement and that the outcome of the circuit – if the single-qubit gates of the circuit are not known, as is the case for Bob as they are redacted for him – is irretrievable from the measurement outcomes without the key. This hiding of the circuit outputs is information-theoretically secure (i.e. has perfect security) [50], assuming the single-qubit gates are redacted, and comparable to the one-time pads in universal blind quantum computing [51]. Henceforth, I will denote $P_{\textit{targ}}$ and $P_{\textit{targ}}$, with the changes implemented by Algorithm 2 to hide their outputs, by $P_{\textit{targ}}^{\prime}$ and $P_{\textit{targ}}^{\prime}$ respectively. I note that if you have the key, the output of the circuit can be easily recovered by XOR-ing each measurement outcome with the corresponding bit in the key.
@@ III.2 Presentation of the Upgraded Accreditation Protocol
With trap and target circuits (with securely encrypted measurement outcomes) established, I can define the full accreditation protocol. In line with Sec. III.1, for Sec. III.2, trap circuits and target circuits will be treated as black boxes and I will only refer to their construction as being performed by the polynomial-time classical algorithms, $P_{\textit{targ}}^{\prime}$ and $P_{\textit{targ}}^{\prime}$. Their only relevant properties will be that trap circuit executions detect any error with probability at least $k\in(0,1]$, all error in a trap or target simulation is effectively reduced to stochastic Pauli error, and Bob cannot tell the difference between trap and target simulations due to them only differing in their single-qubit gates (which makes them indistinguishable to Bob, during Protocol 1, due to the redaction of all single-qubit gates, by Robert, before they are shown to Bob).
However, before the full protocol can be presented, in Protocol 3, I must establish the statistical foundations of the new protocol – in Sec. III.2.1 – and the core mechanics of the protocol – in Sec. III.2.2. These statistical methods will be used to evaluate the $P_{0}\in[0,1]$ that defines Bob’s particular choice of SPSCLβ (as in Def. 12), using multiple trap executions ( all within a single use of Protocol 1). If a target circuit is then executed within that same single use of Protocol 1 as the trap circuit executions, this allows the probability of error (as in Def. 11) of the execution of the target circuit to be bounded upper (using that all circuit executions in a single use of Protocol 1 have error within a single SPSCLβ). Due to the argument in Ref. [23, Appendix Sec. 1], this, in turn, upper bounds the ideal-actual variation distance (as defined in Def. 1) of the target circuit execution.